Find the sum of the first 10 terms of the series 6.4, 3.2, 1.6, ...
Answer by stanbon(48559)

Find the sum of the first 10 terms of the series 6.4, 3.2, 1.6, ...
------------
r = 3/2/6.4 = 1/2
------
S(10) = 6.4[(1/2)^11 -1)/((1/2)-1)]
S(10) = 6.4[-0.9995117/-0.5] = 12.79375
==============
Cheers,
Stan H.